What is the equation of the line perpindicular to y=-1/5x+2 that passes through point (10,7)?
Answer by ewatrrr(24785)


Hi


y = (-1/5)x + 2 |m = (-1/5) Line ⊥ to it: m = 5 (the negative reciprocal)
New Line
y = 5x + b |using ordered pair Pt(10,7) to solve for b
7 = 5*10 + b
-43 = b
y = 5x - 43